# Yuan Luo1, Lizhuang Ma1, Shuicheng Yan2

### v2

@@ 1 Introduction
Humans are easily able to recognize images with domain distribution shifts (such as background changes [4] and various lighting conditions [56]) since the main semantic concepts are consistent. However, this is challenging for multimedia [43, 54, 48, 68, 100, 95, 19, 41, 26, 96] and computer vision systems [53, 15, 3, 27, 28, 93, 94, 59]. One effective approach for eliminating distribution shifts is domain generalization (DG) [45, 89, 20, 5], which attempts to encourage models to focus on semantic factors akin to humans and overlook non-essential features  [21, 72, 81, 87, 17, 89, 70, 38].
A large amount of research in DG has concentrated on the design of special modules to acquire robust representation [81, 98, 65, 99, 97, 77], including domain alignment [16, 84, 30, 46], feature contrastive learning [32, 47], and style augmentation [86, 92, 8, 68, 78], and etc. Nonetheless, prevailing DG methods heavily rely on CNNs as the backbone to extract latent features, only possessing local receptive fields. As a result, they tend to learn local details and overlook global information, impeding generalization and leading to less-desired performances on unseen domains. Recent advancements in DG [35, 50, 88] have shifted the backbone architecture from CNN [25] to ViT [11] due to its global receptive fields of self-attention layer [35, 51, 50, 22]. However, such attention layers in ViT introduce the challenge of quadratic complexity and lead to unacceptable computational inefficiency and memory overhead especially the models are very large, which makes it hard to deploy these DG methods in real-world applications.
@@
Mamba, as an emerging state space model (SSM), possesses superior linear complexity and global receptive fields. It has been recently explored in language modeling and has a promising potential in computer vision. By employing input-dependent parameters in SSM, Mamba has exhibited promising performance in sequence data modeling and capturing long-range dependencies. In particular, VMamba [44] and Vision Mamba [101] propose to traverse the spatial domain and convert any non-causal visual image into order patch sequences. However, such SSM-based models inevitably exhibit performance degradation in unseen domains due to the lack of consideration of domain shifts and tailored designs, and there still exist generalization performance gaps compared with the state-of-the-art DG methods, e.g,, iDAG [30] on the PACS dataset (87.7% vs. 88.8%). Therefore, how to improve the generalizability of Mamba-based models and investigate what hinders Mamba from distribution shifts for DG is a very critical problem.
@@ 2 Related Work
CNN-based models in DG employ convolution neural networks, e.g., Alexnet [33] and ResNet [25], to extract stronger representations. These approaches specially designed submodules to regularize the acquired features based on the CNN backbones. The most intuitive idea for DG is to minimize the empirical source risk [45, 83, 73, 1, 29]. Domain alignment [16, 39, 84, 82, 66] constrained models to convey little domain characteristics by an extra domain discriminative network. To learn domain-invariant representations, feature disentanglement methods [76, 68, 67, 40] aim to disentangle the features to acquire task-specific information. Another significant avenue is to augment the source data [58, 90, 90, 92, 64], thereby providing models with more samples with diverse styles. Inspired by the expressive performance of self-supervised learning, contrastive learning [74, 32, 47, 31, 31, 31] employed the contrastive loss function on features to reduce the gap of representation distributions in one category. Motivated by the efficacy of smooth loss landscapes in enhancing generalization performance, ensemble learning [7, 9, 32] utilized stochastic weight average to find a flatter minimum in loss spaces. Approaches based on meta-learning [85, 12, 13, 2] attempt to address the distribution shift issue during the training phase, enabling models to learn to tackle domain shifts. Despite their remarkable progress in DG and the linear complexity, the lack of global receptive fields hinders further developments of CNN-based models for boosting generalization performance [35, 44].
@@ 3.1 Preliminaries
Shortcomings of Mamba for DG. However, the naive Mamba network encounters challenges when confronted with distribution shifts in DG, achieving a generalization performance of 87.7% on PACS dataset, inferior to the existing DG method iDAG [30] (88.8%). In the following section, we will delve into the reason behind this performance gap and propose effective strategies to assist Mamba in tackling distribution shifts for DG.
@@ 3.3 Semantic-aware Patch Refining
Context Patch identifying. To distinguish the context and object patches, we take advantage of Grad-CAM [57] as the metric to measure the contributions of different regions of images. As the regions containing the object would activate the Grad-CAM greatly, while the patches exhibiting the context possess a low value in the Grad-CAM, we split the image patches into context and object according to their values in the activation map generated by Grad-CAM. Specifically, the patches with the smallest 25% Grad-CAM values are determined as the context information $z_{c}$, while the remaining are used as object information $z_{o}$.
@@ 4 Experiments
DGMamba VMamba-T - 98.9 64.3 79.2 80.8 80.8
Implementation details. Our proposed model employs Mamba [18] as the backbone, which is pretrained on ImageNet [10]. Following VMamba [44], the network comprises 4 blocks, each consisting of 2, 2, 4, and 2 Mamba layers, respectively. Down-sampling is incorporated before each block. Additionally, bidirectional Mamba is utilized to enable each patch to gather information from any other patches. Following the training configuration in existing DG approaches [20, 7, 35], the model undergoes training for 10000 iterations, with a batch size of 16 for each source domain. For optimization, we employ the AdamW optimizer with betas of (0.9, 0.999) and a momentum of 0.9. We incorporate a cosine decay learning rate scheduler. The initial learning rate is searched in [3e-4, 4.5e-4].
@@ 4.2 Ablation study and analysis
iDAG [30] (ICCV’2023) ResNet50 25 8 94 88.8
iDAG [30] (ICCV’2023) ResNet101 43 15 495 89.2
@@
DGMamba (ours) VMamba-T 31 5 233 91.2
@@ 5 Conclusion
This work is the first endeavor to explore the generalizability of the SSM-based model (Mamba) in DG. We propose a novel framework named DGMamba, that contains two pivotal techniques. Firstly, we design a novel Hidden State Suppressing to alleviate the adverse effect of domain-specific information conveyed in hidden states. Secondly, we propose Semantic-aware Patch Refining, which consists of Prior-Free Scanning and Domain Context Interchange. Both aim to direct the model’s focus toward the object rather than the context. Extensive experiments on five widely used DG benchmarks show the superiority of DGMamba compared with SOTA DG methods based on CNN or ViT. We believe our work builds a solid baseline for exploiting SSMs for the DG community.
Future work. In the future, we would like to investigate the feature prompt or domain prompt to facilitate SSM-based models learning more powerful representations for DG. This strategy has demonstrated effectiveness in guiding model learning with ViT. Besides, it is valuable to delve further into Mamba architecture in high-structure tasks to unleash its potential in addressing distribution shifts, such as domain-generalized semantic segmentation and domain-generalized object detection.
@@ A.1 Experiment Details
Hyperparameters for DomainNet. The benchmark DomainNet [52] presents a great challenge, containing 586575 images. Training 10,000 iterations represents less than two complete epochs on DomainNet, which is insufficient for the model to converge effectively. Consequently, recent state-of-the-art DG works [7, 35, 82] increase the iterations to 15000 with a batch size of 32 for each domain. To make a fair comparison while considering the constraints posed by GPU, our proposed DGMamba undergoes 80000 iterations with a batch size of 6 for each source domain. The corresponding initial learning rate is searched in [7e-4, 8.5e-4].
@@ A.2 Additional Experiments
DGMamba (ours) VMamba-T 31M 91.3 87.0 99.0 87.3 91.2

### v3

@@ 1 Introduction
Humans are easily able to recognize images with domain distribution shifts (such as background changes [4] and various lighting conditions [55]) since the main semantic concepts are consistent. However, this is challenging for multimedia [43, 53, 48, 96, 98, 93, 19, 41, 26, 94] and computer vision systems [15, 3, 95, 27, 91, 92, 58]. One effective approach for eliminating distribution shifts is domain generalization (DG) [45, 87, 20, 5], which attempts to encourage models to focus on semantic factors akin to humans and overlook non-essential features  [21, 70, 31, 85, 17, 87, 68, 38].
A large amount of research in DG has concentrated on the design of special modules to acquire robust representation [31, 96, 64, 97, 95, 75, 92], including domain alignment [16, 82, 29, 46], feature contrastive learning [32, 47], and style augmentation [84, 89, 8, 96, 76], and etc. Nonetheless, prevailing DG methods heavily rely on CNNs as the backbone to extract latent features, only possessing local receptive fields. As a result, they tend to learn local details and overlook global information, impeding generalization and leading to less-desired performances on unseen domains. Recent advancements in DG [35, 50, 86] have shifted the backbone architecture from CNN [25] to ViT [11] due to its global receptive fields of self-attention layer [35, 51, 50, 22]. However, such attention layers in ViT introduce the challenge of quadratic complexity and lead to unacceptable computational inefficiency and memory overhead especially the models are very large, which makes it hard to deploy these DG methods in real-world applications.
@@
Mamba [18], as an emerging state space model (SSM), possesses superior linear complexity and global receptive fields. It has been recently explored in language modeling and has a promising potential in computer vision [44, 99]. By employing input-dependent parameters in SSM, Mamba has exhibited promising performance in sequence data modeling and capturing long-range dependencies. In particular, VMamba [44] and Vision Mamba [99] propose to traverse the spatial domain and convert any non-causal visual image into order patch sequences. However, such SSM-based models inevitably exhibit performance degradation in unseen domains due to the lack of consideration of domain shifts and tailored designs, and there still exist generalization performance gaps compared with the state-of-the-art DG methods, e.g,, iDAG [29] on the PACS dataset (87.7% vs. 88.8%). Therefore, how to improve the generalizability of Mamba-based models and investigate what hinders Mamba from distribution shifts for DG is a very critical problem.
@@ 2 Related Work
CNN-based models in DG employ convolution neural networks, e.g., Alexnet [33] and ResNet [25], to extract stronger representations. These approaches specially designed submodules to regularize the acquired features based on the CNN backbones. The most intuitive idea for DG is to minimize the empirical source risk [45, 81, 71, 1, 28]. Domain alignment [16, 39, 82, 65] constrained models to convey little domain characteristics by an extra domain discriminative network. To learn domain-invariant representations, feature disentanglement methods [74, 66, 40] aim to disentangle the features to acquire task-specific information. Another significant avenue is to augment the source data [57, 88, 88, 89, 63], thereby providing models with more samples with diverse styles. Inspired by the expressive performance of self-supervised learning, contrastive learning [72, 32, 47, 30, 30, 30] employed the contrastive loss function on features to reduce the gap of representation distributions in one category. Motivated by the efficacy of smooth loss landscapes in enhancing generalization performance, ensemble learning [7, 9, 32] utilized stochastic weight average to find a flatter minimum in loss spaces. Approaches based on meta-learning [83, 12, 13, 2] attempt to address the distribution shift issue during the training phase, enabling models to learn to tackle domain shifts. Despite their remarkable progress in DG, the lack of global receptive fields hinders further developments of CNN-based models for boosting generalization performance [35, 44].
@@ 3.1 Preliminaries
Shortcomings of Mamba for DG. The naive Mamba network encounters challenges when confronted with distribution shifts in DG, achieving a generalization performance of 87.7% on PACS dataset, inferior to the existing DG method iDAG [29] (88.8%). In the following section, we will delve into the reason behind this performance gap and propose effective strategies to assist Mamba in tackling distribution shifts for DG.
@@ 3.3 Semantic-aware Patch Refining
Context Patch identifying. To distinguish the context and object patches, we take advantage of Grad-CAM [56] as the metric to measure the contributions of different regions of images. We utilize the training model without additional models to obtain the gradients and then evaluate the Grad-CAM. The Grad-CAM evaluation is used only in training. As the regions containing the object activate the Grad-CAM greatly, while the patches exhibiting the context possess a low value in the Grad-CAM, we split the image patches into context and object according to their values in the activation map generated by Grad-CAM. Specifically, the patches with the smallest 25% Grad-CAM values are determined as the context information $z_{c}$, while the remaining are used as object information $z_{o}$.
@@ 4 Experiments
DGMamba VMamba-T - 97.7 64.8 79.3 81.0 80.7
Implementation details. Our proposed model employs Mamba [18] as the backbone, which is pretrained on ImageNet [10]. Following VMamba [44], the network comprises 4 blocks, each consisting of 2, 2, 4, and 2 Mamba layers, respectively. Down-sampling is incorporated before each block. Additionally, bidirectional Mamba is utilized to enable each patch to gather information from any other patches. Following the training configuration in existing DG approaches [20, 7, 35], the model undergoes training for 10000 iterations, with a batch size of 16 for each source domain. For optimization, we employ the AdamW optimizer with betas of (0.9, 0.999) and a momentum of 0.9. We incorporate a cosine decay learning rate scheduler. The initial learning rate is searched in (3e-4, 5e-4).
@@ 4.2 Ablation study and analysis
iDAG [29] (ICCV’2023) ResNet50 23 8 94 88.8
iDAG [29] (ICCV’2023) ResNet101 41 15 495 89.2
@@
DGMamba (ours) VMamba-T 31 5 233 91.0
@@ 5 Conclusion
This work is the first endeavor to explore the generalizability of the SSM-based model (Mamba) in DG. We propose a novel framework named DGMamba, that contains two pivotal techniques. Firstly, we design a novel Hidden State Suppressing to alleviate the adverse effect of domain-specific information conveyed in hidden states. Secondly, we propose Semantic-aware Patch Refining, which consists of Prior-Free Scanning and Domain Context Interchange. Both aim to direct the model’s focus toward the object rather than the context. Extensive experiments on five widely used DG benchmarks show the superiority of DGMamba compared with SOTA DG methods based on CNN or ViT. We believe our work builds a solid baseline for exploiting SSMs for the DG community. In the future, we intend to explore the efficacy of our proposed strategy across a broader range of environments with different SSMs. Besides, we would like to investigate the feature prompt or domain prompt to facilitate SSM-based models learning more powerful representations for enhancing the model generalizability.
@@ A.1 Experiment Details
Hyperparameters for DomainNet. The benchmark DomainNet [52] presents a great challenge, containing 586575 images. Training 10,000 iterations represents less than two complete epochs on DomainNet, which is insufficient for the model to converge effectively. Consequently, recent state-of-the-art DG works [7, 35, 80] increase the iterations to 15000 with a batch size of 32 for each domain. To make a fair comparison while considering the constraints posed by GPU, our proposed DGMamba undergoes 80000 iterations with a batch size of 6 for each source domain. The corresponding initial learning rate is searched in (7e-4, 8.5e-4).
@@ A.2 Additional Experiments
DGMamba (ours) VMamba-T 31M 91.2 86.9 98.9 87.0 91.0